OpenStudy (anonymous):

find the coordinates of missing endpoint given that N is the midpoint of MO. N(1,-2) O(-3,-8)

OpenStudy (amistre64):

+4 +4 <......-3.........1...........( )........> o n m <......-8........-2...........( )........> +6 +6

OpenStudy (amistre64):

M(5,4) if i see it right

myininaya (myininaya):

*----*----* M N O to find the midpoint of MO we do \[(\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\] but we are given the midpoint is (1,-2) so we have \[(\frac{x_1+x_2}{2}=1, \frac{y_1+y_2}{2}=-2)\] but we also know one of the endpoints \[(x_1,y_1)\]=(-3,-8) so we have \[(\frac{-3+x_2}{2}=1, \frac{-8+y_2}{2}=-2)\] now solve for both of these equations and you will get your other missing endpoint

OpenStudy (mathteacher1729):

Count how many units you go down from O and how many units you go Left from O to get to N. Do that again, and you'll be at the point M. (see pic)

OpenStudy (amistre64):

you seem to have gotten lost in your latex lol

OpenStudy (amistre64):

yeah, and you complicated it all ways to sunday ;)

myininaya (myininaya):

(-3+x=2, -8+y=-4) (x=5,y=4) (5,4) is the other endpoint gj amistre

myininaya (myininaya):

i like my way :)

OpenStudy (anonymous):

is the answer (-1,3)

myininaya (myininaya):

(5,4) is the answer